# Knowledge Graph Pruning for Recommendation

### v1

@@ 4.2.1. Overall Performance
In order to demonstrate the superiority of our proposed KGTrimmer methods for knowledge graph recommendation, we compare it with other prune baselines under two state-of-the-art knowledge graph recommendation backbones, and the results are shown in Table 2. From the results, we can get several observations.
(1) The performance of our proposed KGTrimmer surpasses most of baselines with the same pruning ratio under all datasets, demonstrating the effectiveness in preserving the crucial information of the knowledge graph for recommendation task. It is because collaborative filtering signals enable our model to better capture the influence of entities on users, while other baselines do not consider the role of users in the pruning task. And thanks to the dual-view module, our model can retain the filtered collective information and beneficial holistic information. (2) The advantages of our method become more evident as the pruning ratio increases. At a pruning ratio of first-class (90% in Amazon-book, 30% in Last-FM and Alibaba-Fashion), the performance of a few baselines remains comparable with our method, suggesting that they indeed filter some valueless or noisy entities. However, when pruning more triplets, our method significantly outperforms them because pruning solely relies on the KG itself has reached a performance bottleneck, while the advantages of our model in capturing collaborative filtering signals gradually become prominent. (3) Different backbones exhibit varying performance towards the quality of the KG. Specifically, KGIN can achieve better performance compared to KGRec at the same pruning ratio. We attribute this gap to the mask mechanism in KGRec which heightens the reliance on the quality of KG. (4) CAGED and KGist are found to be suboptimal in this task, especially in Amazon-book dataset. Although they perform well in noisy and error detection task, they are not originally designed for identifying valueless node and thus fail to recognize the true but redundant information.
@@ 4.2.2. Comparison Between Entire KG and the Pruned KG
The ultimate goal of pruning is to maintain the prediction accuracy while speeding up our prediction, especially in large-scale knowledge graph recommendation tasks. As a result, we compare the pruned knowledge graph generated from KGTrimmer with original one under the KGIN backbone. From the Table 3, we observe that (1) The performance trained on the pruned KG remains comparable with the whole KG, or even surpasses that on Alibaba-Fashion dataset. This observation verifies our motivation that the presence of numerous valueless nodes and noisy nodes in the KG may harm the recommendation performance. (2) KG pruning can lead to significant time efficiency improvements. It is worth noting that the training efficiency has increased by 12.2 times without any compromise in performance in Amazon-book dataset. Even though the number of user-item interactions is several times higher than the number of triplets in the Last-FM and Alibaba-Fashion dataset, KG pruning still offers substantial time efficiency gains. This phenomenon can be attributed to sophisticated structures designed on the KG in the existing models, which can boost performance but also increase the training time. (3) Different datasets contain varying number of valueless nodes and noisy nodes. To be specific, in the Amazon-book dataset, performance deterioration only occurs after pruning up to 95%, while in Last-FM, performance decline is seen at 30%, and in the Alibaba-Fashion dataset, a relatively good performance is still maintained even after 70% pruning.
4.2.3. Effect of Pruning Threshold
During experimental phase, we employ pruning ratio to ensure a fair comparison among baselines. In reality, it is rational to utilize threshold $\tau$ to prune the KG, namely all triplets below $\tau$ will be removed.
@@ 4.2.4. Ablation Study
In this part, we conduct the ablation study to show how the proposed dual-view importance evaluator affects the final performance of KGTrimmer in Table 4. Specifically, removing the collective view results in the dramatic decline of overall performance, undoubtedly proving the indispensable function of the collective view. On the other hand, the variant without holistic view also exhibits inferior performance , particularly in Amazon-Book dataset, which highlights the importance of capturing the impact of entities on recommendation tasks.
@@ 4.2.5. Parameters Sensitivity
The parameter $\gamma$ is primarily used to balance the weights between the two views. Specifically, when $\gamma=1$, only the impact of the collective view is considered, whereas when $\gamma=0$ , only the impact of the holistic view is taken into account. The results of these two aspects have already been demonstrated in the ablation studies. From Figure 4, we have arrived at the following conclusion : (1) Setting $\gamma=0.5$ services as a favorable trade-off for different datasets, indicating that both collective view and holistic view are beneficial to the task. 2) The Amazon-Book dataset experiences a dramatically decline at $\gamma$ values of 0.1 and 0.9, indicating that it relies heavily on both the collective and holistic views. Meanwhile, Alibaba-Fashion and Last-FM see a sharp drop at a $\gamma$ value of 0.1, demonstrating that the collective view is of crucial importance for them.
5. RELATED WORK
5.1. Knowledge-Aware Recommendation
Path-based models (Catherine and Cohen, 2016; Hu et al., 2018; Ma et al., 2019; Sun et al., 2018; Wang et al., 2018, 2019b) focus on the patterns of the long-connectivity among users and entities via items. These patterns, such as meta-path (Hu et al., 2018; Jin et al., 2020), are delicately designed to capture the user preference on certain entities, and thus shed some light on the interpretation. However, these schemes are gradually fading out due to the heavy dependency on pattern quality.
Embedding-based models (Ai et al., 2018; Cao et al., 2019; Huang et al., 2018; Wang et al., 2020; Zhang et al., 2016) attempt to incorporate the semantic information in knowledge graph. They equip items with the KGE (knowledge graph embedding) techniques to discover latent relationships among entities. Despite achieving impressive performance on recommendation tasks, these models overlook the higher-order connectivity within the graph.
GNN-based models (Wang et al., 2021a, 2019c; Du et al., 2022; Wei et al., 2022) target at aggregating messages from neighbors to enrich the ego node embedding. which inherits the propagation mechanism of GNN(Zhou et al., 2020). Typically, KGIN (Wang et al., 2021a) proposes an end-to-end fashion to learn the latent user intents and provide more insight on interpretability. KGRec(Yang et al., 2023) proposes a rational-aware masking mechanism to identify the most important information in the KG, which mitigate the influence of the noisy and irrelevant triplets. Although they have achieved stunning success, all of them suffer from extremely high computational costs due to the cumbersome KG. Along this line, we propose a simple yet effective method to prune the undesirable nodes and maintain comparable performance on various models and datasets.
5.2. Knowledge Graph Pruning
Furthermore, researchers have devoted themselves to another line to slim the KGs, called Graph Summarization (GS), which compresses the vast KG into a compact one(Liu et al., 2018; Shabani et al., 2023; Doerr and Blenn, 2013; Sacenti et al., 2022; Belth et al., 2020; Kang et al., 2022; Brody et al., 2022; Goyal et al., 2020). One brunch of these approaches leverage machine learning or statistical rules to aggregate nodes. Some works (Doerr and Blenn, 2013; Sacenti et al., 2022) adopt clustering techniques to aggregate neighbors into a super node. Despite the prevalence of these methods, they still are incapable to incorporate the collaborative signals. KGist(Belth et al., 2020) tries to capture the meta-path rules according to the Minimum Description Length principle (MDL). By applying these rules, abnormal edges are identified and eliminated. Another stream of works turns to GNN for help(Kang et al., 2022; Brody et al., 2022; Goyal et al., 2020). For instance, PeGaSus (Kang et al., 2022) optimizes the summarization of large-scale graphs by considering both user preferences and interests, along with the structural properties of the graph. Regardless of remarkable achievements, KG summarization skills still face undesirable limitations. On the one hand, summarized nodes from the aggregated nodes may deviate from the facts that nodes in knowledge graph represent symbolic abstractions of entities from the real-world. On the other hand, they all put emphasis on the role of local neighbors, the signals from the holistic view are greatly neglected. Along this line, we launched an unexplored task knowledge graph pruning for the recommender system and proposed a novel dual-view-based framework to address the two issues above elegantly.

### v2

@@ 5.2.1. Overall Performance
In order to demonstrate the superiority of our proposed KGTrimmer methods for knowledge graph recommendation, we compare it with other prune baselines under two state-of-the-art knowledge graph recommendation backbones, and the results are shown in Table 2 and 3. From the results, we can get several observations.
  • The performance of our proposed KGTrimmer surpasses most of baselines with the same pruning ratio under all datasets, demonstrating the effectiveness in preserving the crucial information of the knowledge graph for recommendation task. It is because collaborative filtering signals enable our model to better capture the influence of entities on users, while other baselines do not consider the role of users in the pruning task. And thanks to the dual-view module, our model can retain the filtered collective information and beneficial holistic information.
  • The advantages of our method become more evident as the pruning ratio increases. At a pruning ratio of first-class (90% in Amazon-book, 30% in Last-FM and Alibaba-Fashion), the performance of a few baselines remains comparable with our method, suggesting that they indeed filter some valueless or noisy entities. However, when pruning more triplets, our method significantly outperforms them because pruning solely relies on the KG itself has reached a performance bottleneck, while the advantages of our model in capturing collaborative filtering signals gradually become prominent.
  • The larger the Knowledge Graph, the more potential valueless nodes it contains. According to Table 1, the Amazon-book dataset includes 2.5 million triples, substantially more than Last-FM and Alibaba-Fashion, which contain merely 460k and 270k triples, respectively. Interestingly, when the dataset is pruned by 90%, backbone performance on the Amazon-book remains relatively unaffected, contrasting with the marked performance degradation observed in the other two datasets. This indicates a prevalence of valueless nodes within the Amazon dataset, while the more compact datasets of Last-FM and Alibaba-Fashion appear to contain a higher proportion of valuable information.
  • Different backbones exhibit varying performance towards the quality of the KG. Specifically, KGIN can achieve better performance compared to KGRec at the same pruning ratio. We attribute this gap to the mask mechanism in KGRec which heightens the reliance on the quality of KG.
  • CAGED, KGist and MoG are found to be suboptimal in this task, especially in Amazon-book dataset. Although they perform well in their corresponding tasks, they are not originally designed for identifying valueless node and thus fail to recognize the redundant information.
  • Although KRDN is also a noise-detecting KG-aware recommendation algorithm, its performance in our task was relatively poor. This is because it focuses on obtaining noise-robust representations rather than carefully considering the graph structure. Therefore, it is not suitable for our pruning task.
@@ 5.2.2. Comparison Between Entire KG and the Pruned KG
The ultimate goal of pruning is to maintain the prediction accuracy while speeding up our prediction, especially in large-scale knowledge graph recommendation tasks. As a result, we compare the pruned knowledge graph generated from KGTrimmer with original one under the KGIN backbone. From the Table 4, we observe that
  • The performance trained on the pruned KG remains comparable with the whole KG, or even surpasses that on Alibaba-Fashion dataset. This observation verifies our motivation that the presence of numerous valueless nodes and noisy nodes in the KG may harm the recommendation performance.
  • KG pruning can lead to significant time efficiency improvements. It is worth noting that the training efficiency has increased by 12.2 times without any compromise in performance in Amazon-book dataset. Even though the number of user-item interactions is several times higher than the number of triplets in the Last-FM and Alibaba-Fashion dataset, KG pruning still offers substantial time efficiency gains. This phenomenon can be attributed to sophisticated structures designed on the KG in the existing models, which can boost performance but also increase the training time.
  • Different datasets contain varying number of valueless nodes and noisy nodes. To be specific, in the Amazon-book dataset, performance deterioration only occurs after pruning up to 95%, while in Last-FM, performance decline is seen at 30%, and in the Alibaba-Fashion dataset, a relatively good performance is still maintained even after 70% pruning.
Table 5. Comparison among various models on time cost of pruning process
Model Amazon-book Last-Fm Alibaba-fashion
Statistical
pop 12.24s 5.12s 3.73s
KGist 1.08min 5.56min 31.10min
Learning-based
MoG 5.98h 50.72min 35.42min
CAGED 4.48h 54.12min 30.36h
GNN-based
KRDN 2.06days 1.71days 1.72days
KGRS 2.00days 13.27h 7.41h
KGTrimmer 1.43days 15.62h 12.05h
5.2.3. Time Cost of Pruning Process
Besides confirming the efficiency of the pruned graph, it is also essential to detail the time complexity of the pruning process. To facilitate clearer distinction among them, we have organized all baselines into three classifications: statistical, learning-based, and GNN-based types. In Table 5, we list time cost of baselines and draw the following observations:
  • The pruning process of statistical methods is very efficient, especially the pop method, which can yield pruning results within one minute. Furthermore, as can be seen in Table 2, the pop method performs well on the Alibaba-Fashion dataset. However, its performance drops significantly on other datasets. Upon deeper investigation, we discovered that the graph structure of the Alibaba-fashion is relatively simple, it primarily consists of outfits and their associated fashion stuffs, such as tops, bottoms, shoes, etc. However, the performance is not as good on more complex graph structures like Last-FM and Amazon-book.
  • Learning-based methods are notably efficient, typically delivering pruning outcomes within half a day. These methods aim to capture the intricate structures of graphs, resulting in impressive performances across a variety of datasets. Nonetheless, their limitation in identifying the significance of collaborative signals leads to a significant performance gap when compared to our pruning approach.
  • GNN-based methods are the most time-consuming, generally taking several days to produce results. This can be attributed to the fact that GNNs require stacking multiple layers to capture higher-order signals, a process that takes more time compared to other methods. However, this is also the reason why GNNs usually have stronger expressive capabilities and superior performance. Compared to other GNN-based methods, our KGTrimmer has roughly the same consumption of processing time, but we particularly emphasize the importance of collaborative signals for the pruning task, thus we can produce a higher-quality pruned graph. Additionally, in the future, we will attempt to replace the importance-aware module with a more efficient framework, thus allowing us to obtain the pruned graph within one day or even one hour.
@@ 5.2.4. Ablation Study
In this part, we conduct the ablation study to show how the proposed dual-view importance evaluator affects the final performance of KGTrimmer. To validate the effectiveness of our dual-view Trimmer, we constructed two model variants. In the first variant, we removed the collective score, denoted as w/o col. The second variant involves removing the holistic view score, denoted as w/o hol. Notably, we conduct experiments on three datasets, employing KGIN as the backbone with a pruning ratio of 98%, 90%, 90% in Amazon-book, Last-Fm, Alibaba-Fashion, repectively, and selecting recall@20 and ndcg@20 as the evaluation metrics.
Based on Table 6, we have made the following discoveries: (1) The model equipped with the dual-view outperformed the other two variants across all datasets, which validates the effectiveness of our overall framework. (2) removing the collective view results in the dramatic decline of overall performance, highlighting the intrinsic importance of the collective view in discerning valueless nodes, which, in turn, substantiates the efficacy of our underlying motivation. (3) the variant without holistic view also exhibits inferior performance , particularly in Amazon-book dataset, which highlights the importance of capturing the impact of entities on recommendation tasks.
@@
5.2.5. Effect of Pruning Threshold
During experimental phase, we employ pruning ratio to ensure a fair comparison among baselines. In reality, it is rational to utilize threshold $\tau$ to prune the KG, namely all triplets below $\tau$ will be removed. More detailed discussion could be found in Section 4.5.
@@ 5.2.6. Parameters Sensitivity
The impact of $\gamma$. The parameter $\gamma$ is primarily used to balance the weights between the two views. Specifically, when $\gamma=1$, only the impact of the collective view is considered, whereas when $\gamma=0$ , only the impact of the holistic view is taken into account. The results of these two aspects have already been demonstrated in the ablation studies. From Figure 4, we have arrived at the following conclusion : (1) Setting $\gamma=0.5$ services as a favorable trade-off for different datasets, indicating that both collective view and holistic view are beneficial to the task. 2) The Amazon-book dataset experiences a dramatically decline at $\gamma$ values of 0.1 and 0.9, indicating that it relies heavily on both the collective and holistic views. Meanwhile, Alibaba-Fashion and Last-FM see a sharp drop at a $\gamma$ value of 0.1, demonstrating that the collective view is of crucial importance for them.
6. Future Work
With the exponential growth of data, Large-scale knowledge graphs emerges for its well-organized structure and human-friendly interpretability. However, they lack effective pruning methods which are paving the way for the widespread application of knowledge graph in various fields, such as recommendation. This paper explores the existence of valueless information within knowledge graphs and makes a first attempt at recognizing these nodes via recommendation signals.
To further approach the ultimate goal, there still exists three challenges need to be addressed: 1) Path-based importance evaluator. In our proposed framework, we focus on acquiring the importance of nodes, while neglecting the potential impact of paths. In fact, we are supposed to comprehensively consider the role played by various information within the paths users take to reach entities to better assess the level of attention these entities receive from users. 2) Efficient importance-aware predictor. Although the GNN-based method proposed in this paper has achieved stunning performance, this module requires substantial computational resources, which hinders the further application in industry filed. Thus, it is urgently needed for efficient implementation of importance-aware predictor. 3) Temporal knowledge graph pruning. Worldwide knowledge is constantly changing, especially in the field of recommendation, where user interests can shift significantly over time. To accommodate swiftly evolving interests, it is essential for the pruning approach to be equipped with the ability to process temporal information. Therefore, effectively integrating temporal knowledge stands as a critical challenge to address in the near future.